OpenStudy (anonymous):

8m^7-10m^5divided by 2m^3

OpenStudy (shamil98):

\[\frac{ 8m^7-10m^5 }{ 2m^3 } = \frac{ 2m^3(4m^4-5m^2) }{ 2m^3 }\] You take the GCF of the numerator and simplify there is one more step, do you know what it that is?

OpenStudy (anonymous):

not sure

OpenStudy (shamil98):

you have 2m^3 in the numerator and denominator this can be canceled out, so what happens then? the equation becomes?